# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v4

@@ Appendix D Dataset
The dataset used for training comprises binary images measuring ($\times 32$ 32) pixels, each sample depicting a straight line going through the image center, starting and ending 222 pixels from the image boundary (leading to 595959 distinct lines of different angles). The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate 300300300 image instances (each line about 555 times). During evaluation, we sample each of the 595959 distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
[…]
To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($\times 32$ 32) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
Additionally, we generate binary images measuring ($\times 64$ 64) pixels by hand that represent more complex structures. These images include line drawings of digits 0-999, characters, and objects such as a house, a star, a wheel, a clock, etc. Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
@@ Appendix E Training Parameters
We train the model with a learning rate of $\eta=0.2$ for 100100100 training cycles (epochs). The model uses $\kappa=10$ competitive neurons with recurrent connections spanning a kernel of $\times 11$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of competitive channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
@@ Appendix F Autoencoder Training
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of 333 and a stride of 222, with progressively increasing channel counts of 323232, 646464, 128128128, and 256256256 channels. The decoder mirrors this structure, using transposed convolutional layers with 256256256, 128128128, 646464, and 323232 channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with 777777777k parameters, utilizing up to 256256256 channels, whereas the CNA model operates with only 444444 channels.
The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 ), and a weight decay of $1\times 10^{-8}$.
To provide a robust comparison, we increase the number of training samples for the autoencoder. Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for 200200200 epochs, using 10,00010,00010 , 000 images per epoch. This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction. The autoencoder is trained with a batch size of 512512512.
@@ Appendix G Quantifying Noise Filtration
[…]
@@
[…]

### v5

@@ Appendix D Dataset
The dataset used for training comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles). The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
To introduce additive Gaussian noise to the afferent input in stage $S2$, we probabilistically flip neurons within each channel with a probability of up to $20\%$. Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location. To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
Additionally, we generate binary images measuring ($64\times 64$) pixels by hand that represent more complex structures. These images include line drawings of digits 0-99, characters, and objects such as a house, a star, a wheel, a clock, etc. Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
@@ Appendix E Training Parameters
We train the model with a learning rate of $\eta=0.2$ for $100$ training cycles (epochs). The model uses $\kappa=10$ competitive neurons with recurrent connections spanning a kernel of $11\times 11$ neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of competitive channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
@@ Appendix F Autoencoder Training
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ channels. The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with 777777k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of $(0.9,0.999)$, and a weight decay of $1\times 10^{-8}$.
To provide a robust comparison, we increase the number of training samples for the autoencoder. Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for $200$ epochs, using $10,000$ images per epoch. This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction. The autoencoder is trained with a batch size of $512$.
@@ Appendix G Quantifying Noise Filtration
The overall efficacy of noise reduction remains consistently high across most parameter configurations, exhibiting a constant noise reduction rate of $\geq\approx 95\%$. This rate persists regardless of the magnitude of the introduced noise. As the noise suppression rate is approximately constant, this means that if there is more noise in the input, there will also be more noise in the output, albeit limited to around 5% of the input noise.
@@
In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$. Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines, which is the case for the CNA but not for the autoencoder (e.g., see Figure 10).